# MITIGATING LONG-TAIL BIAS VIA PROMPT-CONTROLLED DIFFUSION AUGMENTATION

### v1

@@ Abstract
Semantic segmentation of high-resolution remote-sensing imagery is critical for urban mapping and land-cover monitoring, yet training data typically exhibits severe long-tailed pixel imbalance. In LoveDA, this challenge is compounded by an explicit Urban/Rural split with distinct appearance and inconsistent class-frequency statistics across domains. We present a prompt-controlled diffusion augmentation framework that synthesizes paired label–image samples with explicit control of both domain and semantic composition. Stage A uses a domain-aware, masked ratio-conditioned discrete diffusion model to generate layouts that satisfy user-specified class-ratio targets while respecting learned co-occurrence structure. Stage B translates layouts into photorealistic, domain-consistent images using Stable Diffusion with ControlNet guidance. Mixing the resulting ratio and domain controlled synthetic pairs with real data yields consistent improvements across multiple segmentation backbones, with gains concentrated on minority classes and improved Urban and Rural generalization, demonstrating controllable augmentation as a practical mechanism to mitigate long-tail bias in remote-sensing segmentation. Source codes, pretrained Models and synthetic datasets are available at Github
@@ I Introduction
Generative augmentation offers a complementary direction by synthesizing additional labeled data rather than only reshaping the loss. Diffusion models provide strong fidelity and diversity [6, 16], and recent Earth-observation work demonstrates diffusion-based satellite image generation and layout-conditioned synthesis [18, 2]. However, for long-tailed segmentation, realism alone is insufficient, and controllability is essential [7]. If a generator primarily matches the empirical training distribution, rare classes remain rare, and domain gaps may be reinforced [11, 22, 14].
In this work, we propose a prompt-controlled generative augmentation framework for LoveDA that explicitly conditions generation on domain (Urban/Rural) and partial class-ratio targets, enabling targeted synthesis that increases minority-class pixel exposure while preserving domain-consistent remote-sensing realism.
@@ IV Conclusion
We proposed a prompt-controlled diffusion-based data augmentation framework for LoveDA that explicitly addresses the coupled challenges of long-tailed pixel imbalance and Urban/Rural domain shift. By conditioning generation on domain identity and class-ratio targets, our approach enables targeted synthesis of samples that increase effective exposure to underrepresented classes while preserving domain-consistent remote-sensing realism. Experiments show that augmenting the original training set with greedily generated Urban and Rural samples consistently improves segmentation performance, particularly for minority classes, and reduces the domain gap across backbones. While ratio adherence degrades for extrapolative targets far from learned co-occurrence statistics and evaluation is currently limited to LoveDA, the results demonstrate that controllable generative augmentation is a practical and effective tool for mitigating long-tail imbalance in remote-sensing segmentation.

### v3

@@ Abstract
Long-tailed class imbalance remains a fundamental obstacle in semantic segmentation of high-resolution remote-sensing imagery, where dominant classes shape learned representations and rare classes are systematically under-segmented. This challenge becomes more acute in cross-domain settings such as LoveDA, which exhibits an explicit Urban/Rural split with substantial appearance differences and inconsistent class-frequency statistics across domains. We propose a prompt-controlled diffusion augmentation framework that generates paired label-image samples with explicit control over semantic composition and domain, enabling targeted enrichment of underrepresented classes rather than indiscriminate dataset expansion. A domain-aware, masked, ratio-conditioned discrete diffusion model first synthesizes layouts that satisfy class-ratio targets while preserving realistic spatial co-occurrence, and a ControlNet-guided diffusion model then renders photorealistic, domain-consistent images from these layouts. When mixed with real data, the resulting synthetic pairs improve multiple segmentation backbones, especially on minority classes and under domain shift, showing that better downstream segmentation comes from adding the right samples in the right proportions.
@@
Source codes, pretrained Models and synthetic datasets are available at buddhi19.github.io/SyntheticGen.
@@ I Introduction
Generative augmentation offers a complementary alternative by synthesizing additional labeled data rather than only reshaping the loss. Diffusion models provide strong fidelity and diversity [5, 15], and recent Earth observation studies have shown their potential for satellite image generation and layout-conditioned synthesis [17, 2]. For long-tailed segmentation, however, realism alone is insufficient; what matters is controllable generation of the right samples. If synthesis merely follows the empirical training distribution, rare classes remain rare and domain imbalance can persist [6, 10, 22, 13].
To address these limitations, we propose a prompt-controlled generative augmentation framework that explicitly conditions synthesis on domain and class-ratio targets. The framework generates domain-consistent label-image pairs that increase minority class exposure while preserving realistic remote sensing structure, thereby improving downstream segmentation not by indiscriminately expanding the dataset, but by adding the right data in the right proportions.
@@ IV Conclusion
We presented a prompt-controlled diffusion augmentation framework for semantic segmentation under long-tailed class imbalance and domain shift in remote-sensing imagery. By conditioning generation on domain identity and class-ratio targets, the method synthesizes label-image pairs that selectively enrich underrepresented classes while preserving domain-consistent realism. Experiments on LoveDA show consistent gains across segmentation backbones, especially for minority classes, together with reduced Urban/Rural domain gap.
More broadly, this work shows that augmentation is more effective when it is controllable across the factors that drive downstream failure, such as class composition and domain. The key value of the proposed framework is therefore not simply to add more data, but to add the right data in the right proportions. Future work will extend this idea to more datasets, tasks, and stronger forms of controllable generative augmentation.